OpenStudy (anonymous):

What are the solutions of the system? y=x2-2x-1 y=-2

OpenStudy (sgadi):

\[y=x^2-2x-1\] \[y=-2\] \[-2=x^2-2x-1\] \[x^2-2x+1=0\] \[(x-1)^2=0\] \[\Rightarrow x=1\]